OpenStudy (anonymous):

log(x+5)=log(x)+log(5)

OpenStudy (anonymous):

\[x = 5/4\]

OpenStudy (anonymous):

can you explain?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Yep, There is a law of logs that says, log of something timeslog of something else = log(somthing+something else). In your case\[log(x) + log(5)\]is actually equal to \[log(5x)\]because of the law of logs which makes your equation\[log(5+x) = log(5x)\]Now we can take the inverse log (or \(e\)) of both sides. to give\[x +5= 5x\] and solve from there. Does that make sense? (Forgive any spelling discrepancies)

OpenStudy (anonymous):

yes it does.

OpenStudy (anonymous):

I got x=4/5 because you subtract the x making it 5=4x

OpenStudy (anonymous):

never mind I got 5/4